OpenStudy (anonymous):

Write an equation of a line in slope-intercept form with a slope of -4 and passing through the point (0, -3)

OpenStudy (anonymous):

Remember y = mx + b Where m is the slope and b is the y-intercept

OpenStudy (anonymous):

Do you know what the slope is?

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

So that means y = -4x + b. Which means all you have to do is find the y-intercept?

OpenStudy (anonymous):

What do you know about the y-intercept?

OpenStudy (anonymous):

-3=-4(0)+-3

OpenStudy (anonymous):

Correct. y = -4(0) + -3 y = -3 which means b = -3

OpenStudy (anonymous):

So then what is your equation?

OpenStudy (anonymous):

.the equation would be -3=-4(0)+-3